OpenStudy (anonymous):

This is life or death please help? Tina is planning to fence in an area for her dog. She wants to create a rectangular area of 1,600 square feet for her pet and can afford to purchase 160 feet of fence. ? In two or more complete sentences, explain the algebraic model, calculations and reasoning necessary to determine the dimensions of the rectangular area.

OpenStudy (gottennis121):

Here, I gtg but some of others might be able to help. @EclipsedStar @sleepyjess @TheSmartone @undeadknight26 @pooja195 @CausticSyndicalist

OpenStudy (anonymous):

life or death? better get busy the largest area rectangle with fixed perimeter is a square if you can only afford 160 feet of fence, then each side of your square is one quarter of that or 40 feet luckily \(40\times 40=1600\) so you just make it

TheSmartOne (thesmartone):

But @satellite73 already got the question :P

OpenStudy (sleepyjess):

Yup

OpenStudy (anonymous):

not really in the sense that i did not make an "algebraic model" but we can do that too if you like

OpenStudy (anonymous):

after all, it is life of death, so i guess we have some responsibility to save a life

OpenStudy (acxbox22):

lol you can count on satellite

OpenStudy (anonymous):

if you call one side of the fence \(l\) and the width \(w\) then \(2l+2w=160\) solving for \(l\) gives \[l=80-w\]so the area is \[A=w(80-w\] or \[A=80w-w^2\] the maximum of that parabola is at the vertex , which is \(-\frac{b}{2a}\) or in your case \(-\frac{80}{-2}=40\)

OpenStudy (anonymous):

as promised, the square does it just enough now go live long and prosper, or whatever

OpenStudy (anonymous):

Lol, thank you so much for explaining @satellite73

TheSmartOne (thesmartone):

@Zarkon we should teach @satellite73 the Moore method ;)

OpenStudy (zarkon):

that would be a good idea ;)

OpenStudy (anonymous):

@satellite73 Is it too much to ask for help on another question?

TheSmartOne (thesmartone):

@satellite73 The Moore method is where the asker does more and the helper does less work. @Zarkon taught it to you on Thu Jun 21 2012 in this post http://openstudy.com/study#/updates/4fe3d125e4b06e92b8728b5d Always good to refresh your memory :P

OpenStudy (acxbox22):

@TheSmartOne how do you know this stuff? you have only been here for half a year and you can retrieve 3 year old posts? hats off to you

OpenStudy (sleepyjess):

lolololol

TheSmartOne (thesmartone):

I have nothing better to do than look through old posts and find interesting stuff. I even know @satellite73 's motto xDDD

TheSmartOne (thesmartone):

His motto is "cheating is learning"

TheSmartOne (thesmartone):

He did say it :P http://openstudy.com/study#/updates/54470a2fe4b0a8c20fc2a1f9